The slope of the AE curve is .80. What is the multiplier? Everything else the same, by how much does equilibrium aggregate expenditure increase if

a) exports increase from $1.75 trillion to $2.25 trillion. (8 marks)

b) government expenditure on goods and services decrease from $2.0 trillion to $1.8 trillion.

(6 marks)

c) investment increases from $1.2 trillion to $2.3 trillion. (6 marks)


1
Expert's answer
2021-09-13T12:19:19-0400

Slope of AE curve = c = 0.8

Expenditure multiplier = k = 1/(1-c)

k = 1/(1-0.8)

k = 1/0.2

k = 5

(1) Change in exports = ΔX = $2.25 - $1.75 = $0.50 trillion

Change in aggregate expenditure = ΔAE

ΔAE/ΔX = k

ΔAE/0.5 = 5

ΔAE = 5 * 0.5

ΔAE = $2.5 trillion

Aggregate expenditure increases by $2.5 trillion.

(2) Change in government expenditure = ΔG = $1.8 - $2 = -$0.20 trillion

Change in aggregate expenditure = ΔAE

ΔAE/ΔG = k

ΔAE/-0.2 = 5

ΔAE = 5 * -0.2

ΔAE = -$1 trillion

Aggregate expenditure deceases by $1 trillion.

(3) Change in investment = ΔI = $2.3 - $1.2 = $1.1 trillion

Change in aggregate expenditure = ΔAE

ΔAE/ΔI = k

ΔAE/1.1= 5

ΔAE = 5 * 1.1

ΔAE = $5.5 trillion

Aggregate expenditure increases by $5.5 trillion.